OpenStudy (anonymous):

What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250? I know n=8, a1=10, an=781,250. I just don't understand how to find the common ratio if a list of the terms is not given and the problem doesn't say what the common ratio is?

OpenStudy (campbell_st):

then use the formula \[T _{n} = ar^{n-1}\] to find the common ratio \[781250 = 10r^{8-1}\] \[r = \sqrt[7]{78125}\] r = 5 next the sum of a geometric series is \[s _{n} = \frac{a( r^n -1)}{r-1}\] substitute your values of n, a and r to find the sum.